Session: Jun 2007 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (TaoAn at Fri Jun 1 22:56:08 2007) Success :

Urumqi (Zhang hua at Fri Jun 1 23:56:22 2007) Success :
The compressor of C-band doesn't working.we used room temperature receiver. Otherwise no known problems.

Medicina (andrea orlati at Sat Jun 2 00:52:42 2007) Success :
The vsn snd the front label of the second disk pack do not match: MPI-0053 in the front, jb-0002 reported in the log file. A part from this, no known problems.

Onsala ( Michael Lindqvist at Sat Jun 2 05:55:35 2007) Success :
No known problems.

Noto (P. Cassaro at Sun Jun 3 10:04:18 2007) Minor failures : FS
From 22:26 to 22:50 UT scan lost due to FS problems.

Torun (Kaz Borkowski at Mon Jun 4 07:15:55 2007) Success :
Four minutes beginning at 17:31 UT lost due to a problem with control of telescope elevation drive. No other known problems.

Effelsberg (A. Polatidis at Mon Jun 4 08:03:02 2007) Minor failures : RECE
VC05 still not working. Observing system restarted at 2150 UT. There may be loss of data before this.

Westerbork (Tony Foley at Mon Jun 4 08:49:23 2007) Failure : RECO FS
Wb used in phased array mode. Big problems with the Westerbork file server (RAID card and possible also disk damage); we had to replace the server on Friday evening 1st June. The file server had also given errors on Thursday. We recorded from 15:30 to 16:44 (operator error had not enabled 2nd bank while checking the local system). Observations restarted at 20:49UT, but most of the data from 20:49 to 00:30 gave large numbers of data_check errors.

Westerbork (Tony Foley at Mon Jun 4 09:13:32 2007) Minor failures :
Single dish was used. Please change the coordinates in the vexfile accordingly. Supplemental : Scan_check gave large errors in the restarted scan at 21:03. scan checks from 16:44 to 21:03 are looking at the same scan.